# TAET: Two-Stage Adversarial Equalization Training on Long-Tailed Distributions

### v1

@@ 2.1 Long-Tailed Recognition
Long-tail distribution is a prevalent form of data imbalance in training datasets, characterized by a small number of ”head” classes with abundant samples, contrasted with a majority of ”tail” classes that have relatively few samples [48]. Models trained on such imbalanced distributions typically show high confidence in predicting head classes, leading to ”overconfidence” that severely hampers generalization to tail classes. This issue becomes particularly pronounced when evaluating rare categories in real-world applications, as illustrated in Fig. 5. Effectively addressing this challenge remains a significant research hurdle [8].
@@ 6.4 Ablation Study
Contribution of HEL. Fig. 6 shows the impact of each component. BCL notably improves accuracy, as expected. HDL and RCEL are designed to boost weak classes and perform best in combination.
@@ B.1 Necessity and Benefits of the TAET Two-Stage Training
The comparison between the two-stage training method (TAET) and the single-stage approach (HAEL) is illustrated in the figure. TAET employs a two-stage training strategy, where the first 40 epochs are dedicated to optimizing cross-entropy (CE), followed by adversarial loss (HAEL) optimization in subsequent epochs. This methodology enables a dynamic equilibrium between natural accuracy and adversarial robustness as shown in Figure 8.
The left graph compares TAET and HAEL in terms of natural accuracy. During the initial 40 epochs of CE training, TAET primarily concentrates on improving the classification accuracy of natural samples. This phase enables rapid convergence and leads to a significant performance improvement over HAEL, thus establishing a strong baseline for natural performance. After completing the 40th epoch, TAET transitions to HAEL optimization. Although a transient fluctuation in natural accuracy occurs during this transition, the model quickly adapts to the new training objective and ultimately achieves a natural accuracy of approximately 0.8, clearly surpassing HAEL. This outcome demonstrates that TAET’s two-stage training approach provides sufficient room for optimizing adversarial robustness while maintaining high natural accuracy, all while offering a substantial reduction in resource requirements.
The right graph presents the comparison of adversarial robustness under PGD attacks. During the first 40 epochs of CE training, TAET predominantly focuses on optimizing natural accuracy, with minimal attention to adversarial robustness, resulting in nearly zero PGD robustness. However, following the switch to HAEL optimization at epoch 40, TAET’s PGD robustness improves at an accelerated rate, eventually surpassing HAEL in the later stages of training and stabilizing around 0.36. This finding underscores that TAET’s strategy effectively prioritizes natural accuracy optimization before shifting focus to adversarial robustness, achieving a dynamic balance between the two objectives. Additionally, by combining CE and HAEL in a complementary manner, TAET not only significantly enhances the model’s ability to generalize on natural samples but also demonstrates substantial potential for improving adversarial robustness, while significantly reducing the computational burden compared to HAEL.These advantages collectively result in nearly halving the time and memory consumption, as shown in Figure 6.
@@ B.2 Conventional Evaluation Metrics on CIFAR10-LT
The experimental results on CIFAR-10-LT using accuracy and robustness metrics are shown in Tab. 6. From these results, it is evident that our method achieves high robustness while maintaining strong performance in terms of conventional accuracy, demonstrating the best performance across multiple adversarial attack scenarios, including PGD CW, and AA. Additionally, we observe that the MART method exhibits robust performance under the long-tail distribution, but its accuracy on clean samples is relatively lower. This suggests that a promising avenue for further improvement could involve combining the MART method with long-tail recognition approaches to enhance robustness in long-tail settings while maintaining higher accuracy on natural samples.
@@ Appendix C Time Efficiency Analysis
C.1 Single Iteration Computational Complexity Model
The computational cost for a single sample can be defined as:
$\begin{split}T_{\text{step}}&=\underbrace{F}_{\text{Forward}}+\underbrace{B}_{\text{Backward (Clean Sample)}}\\ &\quad+\kappa\cdot\Big{(}\underbrace{F}_{\text{Forward}}+\underbrace{B_{\text{adv}}}_{\text{Backward (Adversarial Sample)}}\Big{)}\end{split}$ (12)
where:
  • $F$: Forward propagation time.
  • $B$: Backward propagation time for clean samples.
  • $B_{\text{adv}}$: Backward propagation time for adversarial samples (including perturbation generation).
  • $\kappa$: Number of adversarial training iterations ($\kappa=10$).
C.2 Total Training Time Comparison
@@
C.3 Acceleration Ratio Derivation
@@
  • $\rho=\frac{F+B}{F+B+\kappa\cdot(F+B_{\text{adv}})}$: Relative efficiency during the CE stage (empirical value: $\rho\approx 0.28$).
  • $\gamma=\frac{F+B_{\text{adv}}}{F+B}$: Adversarial training time expansion coefficient (empirical value: $\gamma\approx 2.5$).
@@
$\eta\approx\frac{40\times 0.28+60\times(1+10\times 2.5)}{100\times(1+10\times 2.5)}=\frac{11.2+1560}{2600}\approx 0.61$ (15)
This indicates that TAET reduces the theoretical computational workload by approximately 39%.
@@ D.1 Memory Requirement Model
  • $M_{\delta}=b\cdot c\cdot h\cdot w\cdot d$: Memory for adversarial perturbation tensors.
@@ D.2 Memory Requirement Comparison
$\Delta M=M_{\text{TAET}}-M_{\text{AT}}=\begin{cases}-M_{\delta}&\text{During the CE stage}\\ 0&\text{During the AT stage}\end{cases}$ (17)
where:
$M_{\delta}=128\times 3\times 32^{2}\times 4\approx 1.5\text{ MB}$ (18)
@@ D.3 Effective Memory Savings
$\xi=\frac{T_{\text{CE}}}{T_{\text{total}}}\cdot M_{\delta}=\frac{40}{100}\times 1.5=0.6\text{ MB}$ (19)
This indicates that TAET reduces the memory occupation by 1.5 MB during 40% of the training time.

### v2

@@ 2.1 Long-Tailed Recognition
Long-tail distribution is a prevalent form of data imbalance in training datasets, characterized by a small number of "head" classes with abundant samples, contrasted with a majority of "tail" classes that have relatively few samples [48]. Models trained on such imbalanced distributions typically show high confidence in predicting head classes, leading to "overconfidence" that severely hampers generalization to tail classes. This issue becomes particularly pronounced when evaluating rare categories in real-world applications, as illustrated in Fig. 5. Effectively addressing this challenge remains a significant research hurdle [8].
@@ 6.4 Ablation Study
Contribution of HARL. Fig. 6 shows the impact of each component. BCL notably improves accuracy, as expected. HDL and RCEL are designed to boost weak classes and perform best in combination.
@@ 7 Conclusion
Acknowledgments
This work was supported in part by the Hefei Municipal Natural Science Foundation under Grant(No.HZR2403), Natural Science Research Project of Colleges and Universities in Anhui Province (No.2022AH051889, No.2308085QF227), the Fundamental Research Funds for the Central Universities and the Fundamental Research Funds for the Central Universities of China.
@@ B.1 Necessity and Benefits of the TAET Two-Stage Training
The comparison between the two-stage training method (TAET) and the single-stage approach (HARL) is illustrated in the figure. TAET employs a two-stage training strategy, where the first 40 epochs are dedicated to optimizing cross-entropy (CE), followed by adversarial loss (HARL) optimization in subsequent epochs. This methodology enables a dynamic equilibrium between natural accuracy and adversarial robustness as shown in Figure 8.
The left graph compares TAET and HARL in terms of natural accuracy. During the initial 40 epochs of CE training, TAET primarily concentrates on improving the classification accuracy of natural samples. This phase enables rapid convergence and leads to a significant performance improvement over HARL, thus establishing a strong baseline for natural performance. After completing the 40th epoch, TAET transitions to HARL optimization. Although a transient fluctuation in natural accuracy occurs during this transition, the model quickly adapts to the new training objective and ultimately achieves a natural accuracy of approximately 0.8, clearly surpassing HARL. This outcome demonstrates that TAET’s two-stage training approach provides sufficient room for optimizing adversarial robustness while maintaining high natural accuracy, all while offering a substantial reduction in resource requirements.
The right graph presents the comparison of adversarial robustness under PGD attacks. During the first 40 epochs of CE training, TAET predominantly focuses on optimizing natural accuracy, with minimal attention to adversarial robustness, resulting in nearly zero PGD robustness. However, following the switch to HARL optimization at epoch 40, TAET’s PGD robustness improves at an accelerated rate, eventually surpassing HARL in the later stages of training and stabilizing around 0.36. This finding underscores that TAET’s strategy effectively prioritizes natural accuracy optimization before shifting focus to adversarial robustness, achieving a dynamic balance between the two objectives. Additionally, by combining CE and HARL in a complementary manner, TAET not only significantly enhances the model’s ability to generalize on natural samples but also demonstrates substantial potential for improving adversarial robustness, while significantly reducing the computational burden compared to HARL. These advantages collectively result in nearly halving the time and memory consumption, as shown in Figure 6.
@@ B.2 Conventional Evaluation Metrics on CIFAR10-LT
The experimental results on CIFAR-10-LT using accuracy and robustness metrics are shown in Tab. 6. From these results, it is evident that our method achieves high robustness while maintaining strong performance in terms of conventional accuracy, demonstrating the best performance across multiple adversarial attack scenarios, including PGD,CW, and AA. Additionally, we observe that the MART method exhibits robust performance under the long-tail distribution, but its accuracy on clean samples is relatively lower. This suggests that a promising avenue for further improvement could involve combining the MART method with long-tail recognition approaches to enhance robustness in long-tail settings while maintaining higher accuracy on natural samples.
@@ Appendix C Time Efficiency Analysis
C.1 Total Training Time Comparison
@@
C.2 Acceleration Ratio Derivation
@@
  • $\rho=\frac{F+B}{F+B+\kappa\cdot(F+B_{\text{adv}})}\approx 0.139$: Relative efficiency during the CE stage
  • $\gamma=\frac{F+B_{\text{adv}}}{F+B}\approx 0.619$: Adversarial training time expansion coefficient
@@
$\eta\approx\frac{40\times 0.139+60\times(1+10\times 0.619)}{100\times(1+10\times 0.619)}\approx 0.608$ (14)
This indicates that TAET reduces the theoretical computational workload by approximately 39.2%
@@ D.1 Memory Requirement Model
  • $M_{\delta}=b\cdot c\cdot h\cdot w\cdot d\approx 1629.42$ MB: Memory for adversarial components
@@ D.2 Memory Requirement Comparison
$\Delta M=M_{\text{TAET}}-M_{\text{AT}}=\begin{cases}-1629.42\text{ MB}&\text{CE stage}\\ 0&\text{AT stage}\end{cases}$ (16)
@@ D.3 Effective Memory Savings
$\xi=\frac{T_{\text{CE}}}{T_{\text{total}}}\cdot M_{\delta}=\frac{40}{100}\times 1629.42\approx 651.77\text{ MB}$ (17)
This indicates that TAET reduces the memory occupation by 1629.42 MB during 40% of the training time.